40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 798     Entry time: Tue Aug 5 10:56:05 2008
Author: Alberto 
Type: Configuration 
Category: General 
Subject: ITMX chamber opened and mirror released 
D-Mass, Steve, Rana, Koji, Yoichi, Alberto,
We opened the ITMX chamber to check the optics after last week earthquake. In particular, from the spectra, ITMX seemed to be stuck and had to be released again. When we inspected the mirror, we found that it wasnít necessary to touch it. It had become free again during the vent thanks to the change of conductivity in the air inside during the vent.
We checked the magnets and they seemed to be fine.
A couple of stop screws had lost the rubber on their tips, although we donít know if that was due to the earthquake.
We also took advantage of the opening to center the LR and the left OSEMs in the mirror to their zero.
Inspecting the table we found a couple of things not totally clear on the configuration of the optics in the table. In particular we found a beam dump located too close to the ifo beam. Eventually we found out that the dump was meant to block a ghost beam coming from the ITM. A better location should probably be figured out for that. We also found that the POXM1 mirror designed to have the maximum reflectivity for the P polarization of the beam at 45 degrees is mounted so that the incident beam is at 22 degrees. This cause the beam to be 90% transmitted and only 10 percent reflected to POX. The transmitted beam appears at ther BSC chamber.

The ifo beam passes so close to the POXM1 mirror so that it can be clipped by its large metal frame ring. The beam at that point is about 6mm large and the ring is about 1cm thick so that we could gain some distance with a different mount.
ELOG V3.1.3-